# A 2-step Framework for Automated Literary Translation Evaluation: Its Promises and Pitfalls

### v1

@@ 2 Background
Building on James S. Holmes’s 1970s’ foundational theories of literary translation, scholars like Susan Bassett, André LeFevere, and Lawrence Venuti have led the “cultural turn” in translation studies, which emphasizes the importance of cultural and historical contexts in literary translation Snell-Hornby (2006). Literary translation has a special place within broader translation studies for its focus on the aesthetic and affective use of language as well as its engagement with fictional content. Some of the persisting concerns in literary translation include the tension between literal (word-for-word) and free or adaptive (sense-for-sense) Munday et al. (2022), the visibility of the translator Venuti (2017), and the social role of the translator Blakesley (2018). More recently, scholars including Kolb and Macken have examined the potential of the integration of technology and machine translation in literary translation processes Kolb (2023); Macken et al. (2022).
@@
String-based metrics like BLEU Papineni et al. (2002) have been widely used in machine translation. Similarly, more recent metrics such as BLEURT Sellam et al. (2020), COMET Rei et al. (2020), and XCOMET Colombo et al. (2023) leverage BERT-style encoder language models Devlin et al. (2019) fine-tuned on translation preference dataset. However, previous work has shown that these metrics are not suitable for literary translation Thai et al. (2022). Departing from earlier trends, Kocmi and Federmann (2023) employed the cutting-edge LLM GPT4 OpenAI (2023) to assess machine translation, where their innovative approach, GEMBA-MQM, set a new benchmark for translation quality estimation at WMT 2023 Freitag et al. (2023). Recent studies Wu et al. (2024) and Chen et al. (2024) also use GPT4 as a judge for literary translation for web novels and Chinese poetry respectively.
We take a step back and conduct a comprehensive study on the LLMs’ capabilities of fine-grained evaluation for literary translation. Another contemporary work, Zhang et al. (2024), has also looked into LLM’s ability to evaluate literary text.
@@ 3 Dataset, Metrics & Annotators
Three literary experts with extensive experience in both English and Korean literature develop and validate the framework presented in Section 4. They also carry out the annotations in Section 5 to validate the framework.
@@ 4 Two-Step Literary Translation Evaluation Framework
DA-MQM: A direct assessment rubric on a scale of 1 to 5, similar to Multidimensional Quality Metric (MQM) Mariana (2014), for evaluating translations using a rubric based on an analysis of language-specific common mistakes found in literary machine translations.
@@ 4.2 Step 2: VERSE
To provide better interpretability for the LLM generated questions in VERSE, we classify the questions into nine categories using an LLM. For all our experiments, the questions are classified in a zero shot manner by GPT4-o. The categories [^3] Note that the categories are not necessarily mutually exclusive and one question might fit into more than one category., as shown in Figure 2, are defined by a human expert after qualitatively examining many evaluation questions. To validate the accuracy of the LLM classification, one human expert manually annotates 150 questions into the nine categories in which GPT4-o scores  77% Top-1 match in accuracy. An examination of the misclassified examples reveals that most misclassifications also fall into a closely related category  [^4] See confusion matrix in Figure 5.
@@ 5 Results
DA-MQM Kendall’s Spear. MSE Krip.
@@
Honorifics .70 .74 .86 .74
Lex. Choices .73 .81 .69 .80
Syntax .66 .76 .86 .75
Content .73 .81 .75 .80
LLM QnA .65 .71 .41 .70
@@
We compare several open-weight and proprietary models, which are known for their good performance in Korean, as evaluators for our experiments. We set the temperature to 0.0 for all experiments in this section for reproducibility.
@@
According to the common practice of literary machine translation Karpinska and Iyyer (2023); Chen et al. (2024); He et al. (2024) , we report COMET Rei et al. (2022), BERTScore Zhang* et al. (2020) and BLEURT Sellam et al. (2020) as the baseline. However, the comparison may not be entirely precise because these metrics are meant to be used in sentence-level evaluation, and our dataset is aligned at paragraph-level.
5.1 DA-MQM: Can LLMs Capture Common Mistakes in Literary Machine Translation?
@@
Honor. 1 30 0.79 0.07 0.81 0.67
2 4 0.15 0.10
3 9 0.11 0.04
4 9 0.22 0.14
5 147 0.91 0.88
Lexical 1 23 0.54 0.43 0.56 0.48
2 29 0.47 0.44
3 42 0.45 0.37
4 50 0.50 0.51
5 55 0.73 0.56
Syntax 1 24 0.61 0.39 0.52 0.44
2 32 0.55 0.44
3 47 0.50 0.27
4 45 0.37 0.43
5 50 0.62 0.58
Content 1 26 0.57 0.19 0.55 0.47
2 29 0.41 0.47
3 42 0.45 0.22
4 49 0.48 0.43
5 53 0.74 0.72
@@
gpt-4o .51 .56 2.03 .68 .77 0.72 .62 .71 0.97 .70 .78 0.81 .63 .70 1.13
claude-3-5-sonnet .40 .41 2.22 .60 .67 1.13 .59 .66 1.17 .57 .64 1.35 .54 .60 1.46
claude-3-5-haiku .42 .45 1.80 .54 .59 1.35 .56 .64 1.22 .55 .62 1.39 .52 .57 1.44
gpt-4o-mini .27 .31 2.74 .63 .70 0.93 .59 .67 0.93 .60 .67 1.08 .52 .59 1.42
claude-3-haiku .22 .23 2.17 .45 .50 1.55 .42 .47 1.47 .51 .56 1.51 .40 .44 1.67
gemma-2-27b-it .34 .38 2.08 .47 .52 1.27 .49 .56 1.24 .49 .55 1.25 .45 .50 1.46
gemma-2-9b-it .28 .30 2.36 .43 .48 1.40 .36 .41 1.47 .52 .58 1.34 .40 .44 1.64
llama-3-8b-it -.01 -.01 2.31 .37 .41 1.56 .29 .32 1.65 .21 .24 2.04 .21 .24 1.89
COMET .10 .12 .28 .35 .25 .32 .28 .36 .23 .29
BERTScore .12 .15 .23 .29 .22 .28 .24 .30 .26 .26
BLEURT .14 .17 .23 .30 .21 .28 .23 .30 .20 .26
@@ 5.2 VERSE: Can LLMs Answer Literary Questions with Subtle Nuance?
Similar to Step 1 DA-MQM, we prompt the language models with human curated rubric and a reference human translation and ask them to give a score.
@@
gpt-4o 0.49 0.53 0.63
claude-3-5-sonnet 0.38 0.41 0.63
claude-3-5-haiku 0.39 0.42 0.66
gpt-4o-mini 0.37 0.40 0.63
claude-3-haiku 0.19 0.20 0.87
gemma-2-27b-it 0.18 0.19 0.70
gemma-2-9b-it 0.09 0.09 0.73
Llama-3.1-8B-it 0.05 0.06 1.06
COMET 0.24 0.31
BERTScore 0.20 0.27
BLEURT 0.20 0.26
@@
Similar to Step 1 DA-MQM, we identify the following key components of LLM’s ability to answer fine-grained questions:
@@
  • Human translation does not help evaluation, in fact not providing human translation is slightly better (Table 13).
@@ 6.1 Experiment Setting
For all the experiments in this section, we use the GPT4-o as the evaluator model for all experiments in zero-shot no CoT setting  [^6] For VERSE, Claude Sonnet 3.5 is slightly better with CoT setting, however, using chain of thought becomes prohibitively expensive for large scale evaluation, which is why we resort to zero shot no CoT GPT4-o, whose performance is close enough. We also did not use reference human translation because in Step 2, reference-free configuration performs slightly better [^7] We also report the evaluation scores computed with reference in the prompt in Appendix H.
@@ 6.2 Fine Grained Literary Evaluation
Step 2 (VERSE) of our framework permits a novel fine-grained interpretation into literary translation, which is otherwise hard to capture through a single number or traditional machine translation metric like fluency, adequacy, etc. Figure 3 shows one such use case of such approach to quickly inspect and visualize different dimensions present in the translation. Moreover, because the language model evaluates  10 questions per paragraph, VERSE estimations are also robust.
@@ 6.3 LLMs prefers LLM generated translation
Table 6 shows the aggregated scores for each Step 1 criteria for each model. Note that since our proposed framework can be run in reference free fashion, we evaluate the reference translations as well.
The results are surprising. Overall, Claude-3.5 Sonnet ($\mathrm{LM}_{G}$) gets the highest scores, followed by GPT4-o, with almost perfect scores in several dimensions. Human translation receives lower scores than the bigger models. To better understand what is going on, we sample 50 paragraphs from the story “What is Remembered", whose human translation receive a particularly lower score than the best performing $\mathrm{LM}_{G}$ translation by ($\mathrm{LM}_{G}$) Claude Sonnet 3.5. In a controlled A/B setting, one human annotator picked the gold translation as the better one 39 out of 50 times, and the machine translation was picked only 4 times.
One potential reason why $\mathrm{LM}_{E}$ is rating human translation lower might have to do with the nature of human translations. Human translations often deviate significantly from their original text in order to properly convey the linguistic nuances. Human inspection reveals that this is indeed the case for the “What is Remembered" story. In fact, for all of our stories, a casual inspection reveals that human translations are clearly better than the best machine translations in majority of the cases.
@@
This suggests that while LLMs might be good at spotting significant differences in translation qualities, they still often fail to distinguish translations with subtle and nuanced differences in Korean. This can be observed from the relatively lower scores provided to Google Translate and GPT3.5, which produce very poor Korean literary translation.
@@
While we keep our analysis limited to English to Korean, the overall structure of evaluation can be used for other languages, provided LLMs have reasonable capability in the target language. Our framework is simple, easy to implement, and provides interpretability suited specifically for literary translation. Much of the pipeline, especially, VERSE is language agnostic and works without much human supervision. For DA-MQM, however, we created the rubric with Korean language in mind, however, this too can easily be adopted with modification for other languages. While gaps remain in current LLMs’ ability to detect some critical issues in literary translation, we believe this paradigm will be helpful for future researchers and practitioners to evaluate long form literary translation.
@@ 7 Conclusion
Our work only tests the framework on one pair of languages, English to Korean. Future works should incorporate other language pairs, which provide different levels of performance. Especially, it remains to be seen how well LLMs can generate literary-specific questions for other non-English languages and then answer them. It is also important to see whether LLMs can take into account different cultural contexts of the source language in the translation. It is possible that LLMs will not be able to generate high-quality questions in low-resource languages or languages that they haven’t been trained on extensively. Similarly, it is also possible that the framework would work even better in higher resource languages.
Acknowledgement
We thank Marzena Karpinska and Sohee Yang for providing some crucial feedback. This work also benefited from fruitful conversations with Graham Neubig and Tom Kocmi.
@@ Appendix A Datast detail
The dataset detail can be found in Table 7.
@@ E.2 Annotation Guidelines for Step 2: LLM Generated Question Answering
    Although it is difficult to ignore awkward expressions or grammatical imperfection in terms of fluency in Korean, always focus on the specific aspect the question asks you to evaluate.
@@ E.3 Annotation Platform
We use Label Studio framework Tkachenko et al. (2020-2022), hosted on Huggingface platform888huggingface.co, to create an annotation platform for all our annotations.
@@ H.2 Detailed Analysis of each criterion of Step 1
We observe some interesting patterns across the score distribution for Step 1 DA-MQM, especially for Korean honorifics.
@@ H.3 Details for of VERSE (Step 2)
Table 8 shows classification statistics averaged over three annotators for each score class. Models tend to do worse when humans gave a Score 1, that is, models struggle to spot a bad translation. This is similar to the behavior observed for honorifics in DA-MQM (Step 1).
@@ H.4 Ablations for DA-MQM (Step 1)
gpt-4o .51 .56 2.03 .68 .77 0.72 .62 .71 0.97 .70 .78 0.81 .63 .70 1.13
w/ CoT .45 .49 1.93 .62 .71 0.86 .60 .69 1.00 .61 .70 1.09 .57 .65 1.22
claude-3-5-sonnet .40 .41 2.22 .60 .67 1.13 .59 .66 1.17 .57 .64 1.35 .54 .60 1.46
w/ CoT .46 .48 2.07 .54 .62 1.17 .56 .64 1.47 .57 .64 1.53 .53 .59 1.56
claude-3-5-haiku .42 .45 1.80 .54 .59 1.35 .56 .64 1.22 .55 .62 1.39 .52 .57 1.44
w/ CoT .45 .48 1.86 .54 .61 1.21 .47 .54 1.58 .35 .39 1.85 .45 .50 1.62
gpt-4o-mini .27 .31 2.74 .63 .70 0.93 .59 .67 0.93 .60 .67 1.08 .52 .59 1.42
w/ CoT .29 .32 3.29 .48 .53 1.53 .48 .55 1.37 .60 .68 1.00 .46 .52 1.80
claude-3-haiku .22 .23 2.17 .45 .50 1.55 .42 .47 1.47 .51 .56 1.51 .40 .44 1.67
w/ CoT .18 .19 2.20 .39 .44 1.71 .43 .48 1.59 .39 .43 1.85 .35 .38 1.84
gemma-2-27b .34 .38 2.08 .47 .52 1.27 .49 .56 1.24 .49 .55 1.25 .45 .50 1.46
w/ CoT .29 .32 2.03 .40 .45 1.44 .37 .42 1.44 .52 .59 1.32 .40 .44 1.56
gemma-2-9b .28 .30 2.36 .43 .48 1.40 .36 .41 1.47 .52 .58 1.34 .40 .44 1.64
w/ CoT .31 .34 2.00 .45 .50 1.35 .32 .37 1.64 .32 .36 1.59 .35 .39 1.64
llama-3-8b -.01 -.01 2.31 .37 .41 1.56 .29 .32 1.65 .21 .24 2.04 .21 .24 1.89
w/ CoT .14 .14 2.44 .15 .16 1.91 .21 .24 1.80 .26 .29 1.81 .19 .21 1.99
@@
As shown in Table 10 providing fine-grained human-curated rubric is important for achieving high correlation with GPT4-o. The correlation and MSE value drops to  0.5 if the detailed rubric is removed. However, different models are affected differently and for some models having a rubric is critical.
@@
gpt-4o .51 .56 2.03 .68 .77 0.72 .62 .71 .97 .70 .78 0.81 .63 .70 1.13
w/o rubric .23 .26 4.63 .68 .77 0.82 .62 .71 0.87 .65 .73 0.85 .54 .62 1.79
w/o human TR .50 .54 1.97 .67 .76 0.83 .63 .72 1.12 .67 .75 1.02 .62 .69 1.24
gpt-4o-mini .27 .31 2.74 .63 .70 0.93 .59 .67 0.93 .60 .67 1.08 .52 .59 1.42
w/o rubric .20 .22 3.34 .64 .71 0.91 .61 .68 0.95 .62 .69 0.97 .51 .57 1.54
w/o human TR .32 .35 2.55 .59 .66 0.96 .56 .63 1.02 .59 .66 1.15 .51 .58 1.42
gemma-2-27b-it .34 .38 2.08 .47 .52 1.27 .49 .56 1.24 .49 .55 1.25 .45 .50 1.46
w/o rubric .32 .33 3.96 .54 .60 1.48 .54 .61 1.17 .59 .66 1.09 .50 .55 1.93
w/o human TR .37 .41 1.91 .52 .58 1.12 .52 .59 1.10 .54 .61 1.17 .49 .55 1.32
gemma-2-9b-it .28 .30 2.36 .43 .48 1.40 .36 .41 1.47 .52 .58 1.34 .40 .44 1.64
w/o rubric .20 .21 3.88 .30 .34 1.83 .29 .33 1.73 .24 .27 1.74 .26 .29 2.29
w/o human TR .29 .32 2.11 .53 .59 1.36 .47 .53 1.52 .50 .56 1.52 .45 .50 1.63
llama-3-8b-it -.01 -.01 2.31 .37 .41 1.56 .29 .32 1.65 .21 .24 2.04 .21 .24 1.89
w/o rubric -.05 -.05 3.20 .18 .19 1.99 .13 .15 2.14 .15 .16 2.18 .10 .12 2.38
w/o human TR .11 .12 2.34 .34 .37 1.65 .29 .32 1.74 .15 .16 2.18 .22 .24 1.98
Table 10: (Step 1) Ablation study results comparing model performance under different conditions. Base model performance is shown in white rows, while ablation conditions (without rubric and without human training) are shown in gray rows. Best results within each model group are in bold. Note that for most models, removing the rubric has a bigger effect than removing the human translation
@@
As seen in Table 10, while human rubrics are good to have, removing them doesn’t lead to a significant performance drop (in terms of correlation) in most criteria. In fact, some models do slightly better without human translation. This is good news since having a human translation for literary text is not always very practical. Moreover, this might also open the door to evaluating different versions of human translations and quantitatively evaluate their differences. However, note that Avg. MSE is higher for w/o human reference cases, which usually means LLMs are biased towards giving higher scores to machine translation when human translation is not provided.

### v2

@@ 2 Background
Building on James S. Holmes’s 1970s’ foundational theories of literary translation, scholars like Susan Bassett, André LeFevere, and Lawrence Venuti have led the “cultural turn” in translation studies, which emphasizes the importance of cultural and historical contexts in literary translation Snell-Hornby (2006). Literary translation has a special place within broader translation studies for its focus on the aesthetic and affective use of language as well as its engagement with fictional content. Some of the persisting concerns in literary translation include the tension between literal (word-for-word) and free or adaptive (sense-for-sense) Munday et al. (2022), the visibility of the translator Venuti (2017), and the social role of the translator Blakesley (2018). More recently, scholars including Kolb and Macken have examined the potential of the integration of technology and machine translation into literary translation processes Kolb (2023); Macken et al. (2022). Our work presents a pipeline incorporating the cultural and aesthetic aspects of literary translation into the evaluation of machine translation.
@@
String-based metrics like BLEU Papineni et al. (2002) have been widely used in machine translation. Similarly, more recent metrics such as BLEURT Sellam et al. (2020), COMET Rei et al. (2020), and XCOMET Colombo et al. (2023) leverage BERT-style encoder language models Devlin et al. (2019) fine-tuned on translation preference dataset. However, previous work has shown that these metrics are not suitable for literary translation Thai et al. (2022). Departing from earlier trends, Kocmi and Federmann (2023) employed the cutting-edge LLM GPT4 OpenAI (2023) to assess machine translation, where their approach, GEMBA-MQM, set a new benchmark for translation quality estimation at WMT 2023 Freitag et al. (2023). Recent studies from Wu et al. (2024) and Chen et al. (2024) also use GPT4 as a judge for literary translation for web novels and Chinese poetry respectively.
We take a step back and conduct a comprehensive study on the LLMs’ capabilities of fine-grained evaluation for literary translation. A contemporary work, Zhang et al. (2024), has also looked into LLM’s ability to evaluate literary text.
@@ 3 Dataset, Metrics & Annotators
Three literary experts with extensive experience in both English and Korean literature develop and validate the framework presented in Section 4. They also carry out the annotations in Section 5 to validate the framework. All annotators were female and native Korean speakers.
@@ 4 Two-Step Literary Translation Evaluation Framework
RULER: RUbric-based Literary Evaluation Rating is a Likert scale of 1 to 5 with four criteria, which resemble Multidimensional Quality Metric (MQM) Mariana (2014), for evaluating translations based on rubrics.
@@ 4.2 Step 2: VERSE
To provide better interpretability for the LLM generated questions in VERSE, we classify the questions into nine categories using an LLM. For all our experiments, the questions are classified in a zero shot manner by GPT4-o. The categories [^3] Notice that the categories are not necessarily mutually exclusive and one question might fit into more than one category., as shown in Figure 2, are defined by a human expert after qualitatively examining many evaluation questions. To validate the accuracy of the LLM classification, one human expert manually annotates 150 questions into the nine categories in which GPT4-o scores  77% Top-1 match in accuracy. An examination of the misclassified examples reveals that most misclassifications also fall into a closely related category  [^4] See confusion matrix in Figure 5.
@@ 5 Results
Criteria Kendall’s Spear. MSE Krip.
@@
Honorifics .70 .74 .89 .73
Lex. Choices .72 .80 .71 .79
Syntax .66 .75 .88 .74
Content .72 .81 .77 .79
LLM QnA .64 .69 .42 .69
@@
We compare several open-weight and proprietary models, which are known for their good performance in Korean, as evaluators for our experiments. We set the temperature to 0.0 for all experiments in this section for reproducibility and report the numbers from a single run from here on.
@@
According to the common practice of literary machine translation Karpinska and Iyyer (2023); Chen et al. (2024); He et al. (2024) , we report COMET Rei et al. (2022), BERTScore Zhang* et al. (2020) and BLEURT Sellam et al. (2020) as the baseline. However, the comparison may not be entirely appropriate because these metrics are meant to be used in sentence-level evaluation, and our dataset is aligned at paragraph-level.
5.1 RULER: Can LLMs Capture Common Mistakes in Literary Machine Translation?
@@
Honor. 1 31 0.79 0.07 0.81 0.66
2 3 0.15 0.10
3 7 0.11 0.04
4 11 0.22 0.14
5 140 0.91 0.88
Lexical 1 16 0.54 0.43 0.54 0.47
2 32 0.47 0.45
3 45 0.45 0.37
4 47 0.50 0.52
5 52 0.70 0.51
Syntax 1 23 0.61 0.39 0.52 0.43
2 32 0.55 0.45
3 49 0.50 0.27
4 43 0.38 0.44
5 45 0.60 0.56
Content 1 19 0.57 0.19 0.54 0.46
2 32 0.41 0.47
3 47 0.45 0.22
4 48 0.47 0.43
5 46 0.73 0.70
@@
gpt-4o .50 .55 2.10 .68 .77 0.71 .62 .71 0.95 .70 .79 0.81 .62 .70 1.14
claude-3-5-sonnet .39 .41 2.30 .58 .65 1.17 .57 .65 1.20 .57 .64 1.38 .53 .59 1.51
claude-3-5-haiku .41 .44 1.86 .54 .60 1.37 .57 .64 1.23 .55 .62 1.42 .52 .58 1.47
gpt-4o-mini .26 .29 2.84 .62 .70 0.93 .60 .68 0.91 .60 .67 1.10 .52 .58 1.44
claude-3-haiku .20 .21 2.24 .45 .50 1.57 .42 .47 1.47 .50 .56 1.55 .39 .43 1.71
gemma-2-27b-it .33 .36 2.16 .46 .51 1.27 .48 .54 1.24 .48 .54 1.26 .44 .49 1.48
gemma-2-9b-it .27 .28 2.44 .42 .47 1.41 .35 .40 1.48 .52 .58 1.35 .39 .43 1.67
llama-3-8b-it -.02 -.02 2.36 .37 .41 1.58 .28 .32 1.67 .21 .23 2.09 .21 .23 1.92
COMET .09 .11 .27 .34 .24 .32 .27 .35 .22 .28
BERTScore .12 .15 .23 .29 .22 .28 .24 .31 .20 .26
BLEURT .13 .16 .22 .28 .20 .26 .22 .29 .19 .25
@@ 5.2 VERSE: Can LLMs Answer Literary Questions with Subtle Nuance?
Similar to Step 1 RULER, we prompt the language models with human curated rubric and a reference human translation and ask them to give a score.
@@
gpt-4o .49 .54 0.61
claude-3-5-sonnet .51 .54 0.73
claude-3-5-haiku .37 .40 0.77
gpt-4o-mini .37 .40 0.61
claude-3-haiku .20 .21 0.82
gemma-2-27b-it .13 .14 0.71
gemma-2-9b-it .09 .09 0.71
Llama-3.1-8B-it .05 .05 1.08
COMET .23 .30
BERTScore .19 .24
BLEURT .19 .24
@@
We identify the following key components of LLM’s ability to answer fine-grained questions:
@@
  • Providing a reference translation does not help evaluation, in fact not providing is slightly better (Table 13).
@@ 6.1 Experiment Setting
For all the experiments in this section, we use the GPT4-o as the evaluator model for all experiments in zero-shot no CoT setting. We also did not use reference human translation because, in Step 2, reference-free configuration performs slightly better [^6] We also report the evaluation scores computed with reference translation in the prompt in Appendix H.
@@ 6.2 Fine Grained Literary Evaluation
VERSE (Step 2) of our framework permits a novel fine-grained interpretation into literary translation, which is otherwise hard to capture through a single number or traditional machine translation metric like fluency, adequacy, etc. Figure 3 shows one such use case of such approach to quickly inspect and visualize different dimensions present in the translation. Moreover, because the language model evaluates  10 questions per paragraph, VERSE estimations are also robust.
@@ 6.3 LLMs prefers LLM generated translation
Table 6 shows the aggregated scores for each Step 1 criteria for each model (see Table 16 for Step 2). Note that since our proposed framework can be run in reference free fashion, we also evaluate the given reference human translations with our framework.
The results are surprising. Overall, Claude-3.5 Sonnet, as $\mathrm{LM}_{G}$, gets the highest scores, followed by GPT4-o, with almost perfect scores in several dimensions. Human translation receives lower scores than the bigger models. To better understand what is going on, we sample 50 paragraphs from the story “What is Remembered", whose human translation receive a particularly lower score than the best performing $\mathrm{LM}_{G}$ translation by ($\mathrm{LM}_{G}$) Claude Sonnet 3.5 (See Figure 8). In a controlled A/B setting, one human annotator picked the gold translation as the better one 39 out of 50 times, and the Claude Sonnet 3.5 translation was picked only 4 times.
An inherent characteristic of human translation may explain why $\mathrm{LM}_{E}$ gives human translation lower scores. Human translations often deviate significantly from the surface-level linguistic features of their original text in order to properly convey the essence of the text. Human inspection reveals that this is indeed the case for the story examined above. In the cases of other stories, too, human translations often prove to read more naturally than the best machine translations in most cases, even when the best machine translations were sufficiently accurate.
@@
This suggests that while LLMs may be good at spotting significant differences in translation qualities, they still often fail to distinguish translations where the difference is more subtle and nuanced. This can be observed from the relatively lower scores provided to Google Translate and GPT3.5, which usually produce very poor Korean translations.
@@
While we keep our analysis limited to English to Korean, the overall structure of evaluation can be used for other languages, provided LLMs have a reasonable understanding of the target language. Our framework is simple, easy to implement, and provides interpretability suited specifically for literary translation. Much of the pipeline, especially, VERSE is language agnostic and works without much human supervision. For RULER, we created the rubric with the Korean language in mind, however, RULER can easily be adopted with modification for other languages. While gaps remain in current LLMs’ ability to detect some critical issues in literary translation, we believe this paradigm will be helpful for future researchers and practitioners to evaluate literary translation.
@@ 7 Conclusion
Our work only tests the framework on one pair of languages, English to Korean. Future works should incorporate other language pairs, which provide different levels of performance. Especially, it remains to be seen how well LLMs can generate literary-specific questions for other non-English languages and then answer them. It is also important to see whether LLMs can take into account different cultural contexts of the source language in the translation. It is possible that LLMs will not be able to evaluate high-quality questions in low-resource languages or languages that they haven’t been trained on extensively. Similarly, it is also possible that the framework works even better in higher-resource languages.
@@ Appendix A Datast detail
Our selected collection of short stories spans nearly two centuries, offering a survey of diverse literary movements and genres in the English language. From classic nineteenth-century gothic and popular suspense stories like Edgar Allan Poe’s “The Masque of the Red Death” (1842) and Charles Dickens’s “The Signal-Man” (1866), to modernist text like Virginia Woolf’s “The Legacy” (1944) and contemporary realist works such as Jhumpa Lahiri’s “Hell-Heaven” (2004) and Andrew Porter’s “Austin” (2017). The genre of speculative fiction tradition is represented by Ursula K. Le Guin’s “The Ones Who Walk Away from Omelas” (1973), Ted Chiang’s “Tower of Babylon” (1990), and Octavia Butler’s “Amnesty” (2003), all of which differ in writing style, while works like Charles Chesnutt’s “The Passing of Grandison” (1899), Katherine Mansfield’s “The Garden Party”(1922) and Tobias Wolff’s "The Deposition" (2006) represent works in which historical and cultural detail contribute greatly to their style. As a collection, these stories cover a wide-range of evolving narrative styles and techniques in English literature.
@@
The names of the stories and other details can be found in Table 7.
@@ E.2 Annotation Guidelines for Step 2: LLM Generated Question Answering
    Although it is difficult to ignore awkward expressions or grammatical imperfections in terms of fluency in Korean, always focus on the specific aspect the question asks you to evaluate.
@@ E.3 Annotation Platform
We use Label Studio framework Tkachenko et al. (2020-2022), hosted on Huggingface platform777huggingface.co, to create an annotation platform for all our annotations.
E.4 Annotated Samples
After the annotation, and some post-processing to discard some samples, we had 188 data points for RULER (Step 1) and 543 data points for validating VERSE (Step 2).
We will release the human annotation for future researchers.
@@ H.2 Detailed Analysis of each criterion of Step 1
We observe some interesting patterns across the score distribution for Step 1 RULER, especially for Korean honorifics.
@@ H.3 Details for of VERSE (Step 2)
Table 8 shows classification statistics averaged over three annotators for each score class. Models tend to do worse when humans gave a Score 1, that is, models struggle to spot a bad translation. This is similar to the behavior observed for honorifics in RULER (Step 1).
@@ H.4 Ablations for RULER (Step 1)
gpt-4o .50 .55 2.10 .68 .77 0.71 .62 .71 0.95 .70 .79 0.81 .62 .70 1.14
w/ CoT .44 .48 2.00 .62 .71 0.86 .59 .68 1.02 .61 .70 1.12 .56 .64 1.25
claude-3-5-sonnet .39 .41 2.30 .58 .65 1.17 .57 .65 1.20 .57 .64 1.38 .53 .59 1.51
w/ CoT .45 .48 2.15 .53 .61 1.21 .55 .63 1.51 .56 .63 1.58 .52 .59 1.61
claude-3-5-haiku .41 .44 1.86 .54 .60 1.37 .57 .64 1.23 .55 .62 1.42 .52 .58 1.47
w/ CoT .44 .47 1.92 .53 .60 1.24 .46 .53 1.62 .32 .36 1.91 .44 .49 1.67
gpt-4o-mini .26 .29 2.84 .62 .70 0.93 .60 .68 0.91 .60 .67 1.10 .52 .58 1.44
w/ CoT .28 .31 3.36 .46 .51 1.51 .47 .54 1.36 .60 .69 1.00 .45 .51 1.81
claude-3-haiku .20 .21 2.24 .45 .50 1.57 .42 .47 1.47 .50 .56 1.55 .39 .43 1.71
w/ CoT .16 .17 2.28 .40 .44 1.74 .44 .49 1.61 .39 .43 1.89 .34 .38 1.88
gemma-2-27b .33 .36 2.16 .46 .51 1.27 .48 .54 1.24 .48 .54 1.26 .44 .49 1.48
w/ CoT .28 .31 2.11 .37 .42 1.46 .35 .40 1.45 .52 .58 1.34 .38 .43 1.59
gemma-2-9b .27 .28 2.44 .42 .47 1.41 .35 .40 1.48 .52 .58 1.35 .39 .43 1.67
w/ CoT .32 .35 2.01 .44 .49 1.35 .29 .33 1.65 .31 .35 1.59 .34 .38 1.65
llama-3-8b -.02 -.02 2.36 .37 .41 1.58 .28 .32 1.67 .21 .23 2.09 .21 .23 1.92
w/ CoT .12 .13 2.50 .13 .14 1.90 .20 .23 1.82 .26 .29 1.84 .18 .20 2.01
@@
As shown in Table 10 providing fine-grained human-curated rubric is important for achieving high correlation with GPT4-o. If the detailed rubric is removed. However, different models are affected differently and for some models having a rubric is critical.
@@
gpt-4o .50 .55 2.10 .68 .77 0.71 .62 .71 0.95 .70 .79 0.81 .62 .70 1.14
w/o rubric .22 .26 4.67 .68 .76 0.81 .62 .71 0.88 .65 .73 0.85 .54 .61 1.80
w/o human TR .50 .53 2.04 .68 .77 0.82 .65 .75 1.10 .69 .77 1.01 .63 .71 1.24
gpt-4o-mini .26 .29 2.84 .62 .70 0.93 .60 .68 0.91 .60 .67 1.10 .52 .58 1.44
w/o rubric .20 .22 3.36 .63 .70 0.92 .62 .69 0.93 .62 .69 0.96 .52 .58 1.54
w/o human TR .30 .34 2.64 .61 .68 0.94 .58 .66 0.99 .60 .68 1.17 .52 .59 1.43
gemma-2-27b-it .33 .36 2.16 .46 .51 1.27 .48 .54 1.24 .48 .54 1.26 .44 .49 1.48
w/o rubric .31 .33 3.96 .51 .57 1.50 .53 .60 1.17 .58 .66 1.10 .48 .54 1.93
w/o human TR .37 .40 1.96 .52 .58 1.12 .52 .59 1.09 .54 .62 1.17 .49 .55 1.34
gemma-2-9b-it .27 .28 2.44 .42 .47 1.41 .35 .40 1.48 .52 .58 1.35 .39 .43 1.67
w/o rubric .20 .21 3.87 .27 .30 1.81 .26 .29 1.72 .20 .22 1.77 .23 .25 2.29
w/o human TR .29 .32 2.15 .54 .60 1.37 .49 .54 1.53 .51 .57 1.54 .46 .51 1.65
llama-3-8b-it -.02 -.02 2.36 .37 .41 1.58 .28 .32 1.67 .21 .23 2.09 .21 .23 1.92
w/o rubric -.04 -.05 3.21 .17 .19 2.03 .13 .15 2.19 .15 .17 2.23 .10 .11 2.42
w/o human TR .11 .12 2.39 .36 .40 1.65 .31 .35 1.74 .15 .17 2.23 .23 .26 2.00
Table 10: (Step 1) Ablation study results comparing model performance under different conditions. Base model performance is shown in white rows, while ablation conditions (without rubric and without human training) are shown in gray rows. Note that for most models, removing the rubric has a bigger effect than removing the human translation
@@
As seen in Table 10, while human rubrics are good to have, removing them doesn’t lead to a significant performance drop (in terms of correlation) in most criteria. In fact, some models do slightly better without human translation. This is good news since having a human translation for literary text is not possible for most of cases. This might also open the door to evaluating different versions of human translations and quantitatively evaluate their differences. However, note that Avg. MSE is higher for w/o human reference cases, which usually means LLMs are biased towards giving higher scores to machine translation when human translation is not provided.